# Forecasting Arctic Temperatures with Temporally Dependent Data Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

### 04 Mar 2026 v3

@@ 5.2 Time-Series Plot for Svalbard Temperatures in 2022, 2023, and 2024
Recall the premise that each time series of daily 2 p.m. temperatures for 2022, 2023, and 2024 represents realizations from the same underlying joint probability distribution; the underlying physics for corresponding months in those three years should be comparable. However, there could be aberrations caused by rapid Arctic amplification.
@@
Figure 2 shows substantial overlap across the three years. Seasonal trends are captured similarly, both in temperature values and in timing. The time series move largely in lockstep from early to late summer, during which melting temperatures are nearly universal. There appears to be no compelling reason to reject the claim that each series consists of random variables drawn from the same joint probability distribution. In practice, the 2022 and 2024 temperatures serve as promising holdout samples for evaluating forecasts of the 2023 temperatures.
@@ 5.3 The Quantile GBM Fit of Temperature
With the quantile parameter fixed at $\tau=0.60$, the quantile gradient-boosting procedure in R ran efficiently. The shrinkage value was set to 0.0001, the interaction depth to 6, and the minimum number of observations in a terminal node to 6. These values were chosen to foster slow convergence so that the right tail of the temperature distribution would be fitted more accurately. Using 2022 as a holdout sample to protect against overfitting indicated that approximately 27,000 iterations were appropriate for these data. [^6] With a substantially larger shrinkage value, many fewer iterations might have been adequate. However, the best shrinkage value could not be known in advance, and empirical tuning encourages “cherry picking” that can undermine later statistical inference (Kuchibhotla, Kolassa and Kuffner, 2022).
@@
Figure 3 plots the observed 2 p.m. temperatures against their fitted values. The relationship is roughly linear and positive; observed and fitted values tend to increase together. Clustering around the smoothed fitted line is somewhat tighter when the fitted values exceed the melting point, with residuals rarely larger than about $\pm 2^{\circ}\mathrm{C}$. This results in part from the larger cost for underfitting imposed by the target quantile $\tau=0.60$. There is greater sparsity on either side of the temperature range between approximately $5^{\circ}\mathrm{C}$ and $0^{\circ}\mathrm{C}$. In short, the fit quality produced by quantile gradient boosting is encouraging. [^7] Koenker and Machado (1999) discuss measures of fit for conditional-quantile models derived from the quantile loss function. However, apparently no analogue of the multiple correlation coefficient exists for quantile regression. A suitable measure of fit is introduced later, when conformal prediction regions are presented.
@@
Overall, the fitted values track the observed temperatures rather closely and capture temperatures above $0^{\circ}\mathrm{C}$ especially well. The unusually warm days clustering in mid-July—around $10^{\circ}\mathrm{C}$—are of particular concern to climate scientists (Semenov, 2021), and they too are fitted accurately, albeit with slight underestimation. Two weeks in advance, the fitted algorithm anticipates melting temperatures effectively, including those mid-July extremes. The fit might improve further if a larger value of $\tau$ were used, but higher quantiles (e.g., $Q(0.90)$) would rely on far sparser data, potentially leading to instability. As with Figure 3, however, true forecasting remains to be performed.
@@ 5.5 Functional Forms
The quantile gradient-boosting algorithm learns associations between each predictor and the response, including the shapes of those relationships. The partial-dependence plots in Figure 6 show the relationships between the day counter in the left and the 14-day lagged temperature on the right with the 2 p.m. temperature response, each computed with all other predictors fixed at their means. These are the two variables that dominate the fit. The black dots represent the partial-dependence values, and the solid blue lines show a loess smooth as a visual aid.
Figure 6: Partial-dependence plots for the day counter and the 14-day lagged 2 p.m. temperature. The black dots are the partial-dependence values, and the solid blue line is a loess smooth serving as a visual aid. Other predictors are fixed at their means.
Consistent with Figure 5, strong relationships are evident. A look again at Figure 4 provides context. The counter-based fitted values range from approximately $3^{\circ}\mathrm{C}$ to $3^{\circ}\mathrm{C}$, and the lagged-temperature fitted values from about $2.5^{\circ}\mathrm{C}$ to $5^{\circ}\mathrm{C}$. The first plot is roughly symmetric and concave, peaking during the summer months, consistent with the longitudinal pattern shown in Figure 4. The second plot is S-shaped, with its steepest slope beginning immediately after the melting point at $0^{\circ}\mathrm{C}$ and plateauing near $2.5^{\circ}\mathrm{C}$ at the left and $5^{\circ}\mathrm{C}$ at the right. The lagged observed temperatures have their strongest association with fitted temperatures just above the melting point. The practical significance of this threshold will be discussed in greater depth below. [^9] The short horizontal strings of dots visible in the plot represent rounding artifacts; the values differ slightly. The remaining four predictors also exhibit nonlinear relationships with the temperature response, but because of their comparatively small contributions to the fit, their partial-dependence plots are omitted in the interest of space.
@@ 6 Forecasts and Estimated Uncertainty
Addressing uncertainty in a principled way has been anticipated in the research design and algorithmic procedures employed. The 2024 data serve as a pristine holdout sample of test data that can represent unlabeled cases used in genuine forecasting. [^10] The 2024 data have labels that are useful for didactic purposes and for evaluation of forecasting accuracy. But those labels have no role in the forecasts themselves and would not be known in an operational setting. Forecasts for the 2024 data can be produced using the previously fitted quantile gradient-boosting algorithm and its predict() function in R. From the perspective of the trained algorithm, the 2024 lagged predictors constitute new, unlabeled data.
Obtaining the adaptive conformal prediction regions is more involved. All conformal prediction regions begin with a pre-specified coverage probability. By convention, this probability is denoted by $(1-\alpha)$, where $\alpha$ lies between 0 and 0.50. In practice, coverage probabilities range from slightly above 0.50 to nearly 1.0, with values between 0.75 and 0.95 most common. The value of $\alpha$ is determined by subjective considerations shaped by the data and the application. For the conformal approach used with the Svalbard data, the coverage is $1-\alpha=0.80$.
@@
Exchangeability is required for valid prediction regions (Vovk, Gammerman and Shafer, 2005; Vovk et al., 2017; Angelopoulos, Barber and Bates, 2024). One implication of exchangeability is that it does not matter, for the theoretical guarantees, that a test observation is realized after all of the calibration observations. A formal summary of conformal prediction is provided in Appendix A along with some computation details. Angelopoulos and Bates (2023) provide an accessible introduction to conformal prediction.
Figure 7 illustrates how adaptive conformal prediction regions work. The figure highlights important features of the Svalbard data but is not meant to be a reproduction of those data. It is a didactic device only.
On the horizontal axis is the forecasted temperature. On the vertical axis is the observed future temperature that would, in practice, be unknown when a forecast is made. The solid blue line conveys how the forecasts and future temperatures are systematically related. The dashed red lines are the upper and lower bounds of the adaptive prediction region for each forecast. If $1-\alpha=0.80$, the prediction region for any forecast contains the correct future temperature with a probability of at least 0.80.
Figure 7: A didactive simulation of adaptive conformal prediction regions from quantile gradient boosting with $\tau=0.80$. The horizontal axis is the forecasted temperature. The vertical axis is the observed future temperature. The solid blue line shows how forecasts and true future temperatures are systematically related. The dashed red lines show the upper and lower bounds of the adaptive prediction region, and the two vertical dotted magenta lines mark the prediction regions when the forecasted temperature happens to be $10^{\circ}\mathrm{C}$ or $5^{\circ}\mathrm{C}$.
To help fix these ideas, consider two illustrative forecasts, one at $10^{\circ}\mathrm{C}$ and the other at $5^{\circ}\mathrm{C}$. Several features of their prediction regions are apparent and apply to all forecasted temperatures in the figure. Forecasts are not necessarily in the middle of their prediction regions, which means that the regions need not be symmetric around their forecasts. In addition, because for the quantile gradient-boosting algorithm $\tau$ was set to 0.60, the forecasts tend to fall above each region’s median. Finally, the lengths of the adaptive prediction regions grow shorter as the fit improves (Romano, Patterson and Candès, 2019). In Figure 7, forecasts above $0^{\circ}\mathrm{C}$ fit the observed temperatures better and have shorter region lengths. In the Svalbard setting, forecasts would be produced one day at a time, but Figure 7 is meant to convey the results after many forecasts are made. The figure also makes plain that there is no requirement that the forecasts be correct.
Although the nominal tail probabilities defining the lower and upper limits are symmetric (e.g., $\alpha/2$ and $1-\alpha/2$), the resulting prediction regions need not be. The asymmetry arises because the values of the upper and lower bounds are determined by the empirical distribution of the nonconformal scores themselves. Their ranks determine the quantiles, but the nonconformal scores need not be equally spaced from one another. For example, if the large positive scores are less dispersed than the large negative scores, the prediction region above the forecasts will be shorter than the prediction region below the forecast.
@@ 6.1 Uncertainty Results
Figure 8: For the 2024 holdout test sample, the mean forecasted temperature in a bin is on the horizontal axis. The percentage of cases in each of 20 equal-width bins having a true future temperature exceeding $0^{\circ}\mathrm{C}$ is on the vertical axis. Filled circles correspond to bins with mean forecasted temperature above $0^{\circ}\mathrm{C}$. Open circles correspond to those below. The solid blue line is a loess smoother.
Figure 8 provides a visual summary of these ideas using the 2024 holdout test data. The horizontal axis shows the mean of the daily forecasted temperatures for each of 20 equal-width bins. Because the future temperatures are known for the 2024 holdout cases, the vertical axis shows the percentage of forecasts within each bin whose true future temperature exceeds $0^{\circ}\mathrm{C}$. (The 14-day lag of each predictor ensures that these are future outcomes.) The loess overlay is S-shaped, with the greatest rate of change between about $2^{\circ}\mathrm{C}$ and $2^{\circ}\mathrm{C}$.
@@
But physical interpretations can be subtle, especially for sea ice. When sea ice forms, its hexagonal lattice cannot readily incorporate salt ions. The excluded salt forms brine pockets and channels among the ice grains, leaving the ice as nearly fresh water while the brine becomes increasingly salty. As freezing continues, the concentrated brine’s freezing point can fall well below $0^{\circ}\mathrm{C}$ because the salt ions attract water molecules the lattice needs; it is more difficult for the water molecules to self-organize into the lattice structure. Unlike fresh water ice, which melts at a single fixed temperature, sea ice melts over a range of temperatures, because the brine pockets within it are already liquid at temperatures below $0^{\circ}\mathrm{C}$. As the temperature rises, progressively more of the remaining ice melts, and the brine channels widen and connect. Because a significant fraction of sea ice consists of these pre-existing liquid brine inclusions, it requires less energy to melt than an equivalent mass of fresh water ice — the brine pockets need no latent heat to liquify, only the nearly pure ice fraction does. By the same logic, the energy released when salt water freezes is less than that released when fresh water freezes. The amount of energy required to melt sea ice equals the amount released when it freezes, consistent with thermodynamic reversibility. This interpretation aligns with Figure 8.
@@ 7 Discussion
The analysis of 2 p.m. temperatures in Svalbard, Norway, combines three interlocking statistical traditions: gradient boosting with a quantile as the estimation target, an AR(1) model to construct exchangeable nonconformal scores, and adaptive conformal prediction regions computed with quantile random forests, which is robust to sparse data. To the best of the author’s knowledge, these procedures applied to three temporally separated samples has not previously appeared in the literature.
A two-week forecasting lead time before the onset of widespread surface melting is a very short fuse, but it would give Arctic community administrators critical time to prepare for infrastructure resilience, public safety, and logistical challenges (Streletskiy, Shiklomanov and Christiansen, 2019; U.S. Arctic Research Commission, 2003). Local governments could limit heavy-vehicle use on thaw-sensitive roads, runways, and around pipelines; pre-position maintenance materials; and inspect culverts, bridges, and drainage channels before damage occurs (Arctic Research Consortium of the United States (ARCUS), 2021). Water managers could adjust reservoir levels or activate temporary treatment measures to mitigate siltation and contamination from meltwater inflows.
In communities dependent on ice roads, snowmobile trails, or frozen river crossings, early warnings would permit orderly resupply and fuel delivery before surface travel becomes unsafe (Chen et al., 2025). Coastal and hillside settlements could prepare for increased risks of shoreline erosion or permafrost-related landslides (The Climate Institute and The Firelight Group, 2022). Health authorities might use the lead time to issue advisories on water quality or vector-borne disease risk as standing water accumulates. Because of Arctic warming, even Iceland is now reported to host mosquitoes (Straker, 2025).
More broadly, forecasts on a two-week timescale could foster coordination among local administrators, regional governments, and research stations, improving situational awareness and resource allocation across sectors. Even modest improvements in lead time may yield substantial adaptive benefits in Arctic regions where melt conditions evolve rapidly and logistical flexibility is limited. [^14] The last three paragraphs were informed by a literature search undertaken by ChatGPT, which also provided some initial wording later edited for consistency and style.
@@ 8 Conclusions
Projections of future global warming have long been central to IPCC assessments, complemented by many smaller-scale forecasting efforts. The analyses presented here contribute to the latter. Temperature forecasts with a two-week lead time appear promising, especially when focused on Arctic thresholds associated with widespread melting. Forecasting uncertainty is explicitly addressed, and results are presented in a manner intended to be accessible and useful to stakeholders. With thoughtfully applied algorithms and accessible data, informative short-term forecasts can be generated on an ordinary desktop computer with an internet connection.
Appendix A. Statistical Narrative
@@
These linked procedures combine quantile gradient boosting, AR(1) whitening, and quantile-random-forest calibration to yield adaptive conformal prediction regions. The lower and upper bounds use the ranks of the nonconformal scores, while the magnitudes of those scores determine the lengths of the resulting prediction regions.
@@
The supervised learning algorithm is first fitted on the training data $\{(X_{t},Y_{t})\}_{t=1}^{T_{\mathrm{train}}}$. Predictions are then generated for the separate calibration data $\{X_{t}\}_{t=1}^{T}$, and the resulting residuals or their AR(1) innovations form the nonconformal scores $\{S_{t}\}_{t=1}^{T}$.
If $\{S_{t}\}_{t=1}^{T+1}$ are (approximately) exchangeable—perhaps after whitening— then, conditional on the fitted algorithm, the calibration scores and the new score $S_{T+1}$ are also exchangeable. Consequently,
@@
Exchangeability means that the joint distribution of the nonconformal scores is invariant to permutations of their indices. Although in practice the scores are realized in temporal order, their joint law would be unchanged had they been realized in any other order. That a new unlabeled case is realized at time T+1T{+}1 is immaterial. For the analyses of the Svalbard data, exchangeability depends on the whitening of the calibration data residuals. The residuals from the training data are not exchangeable, and neither are the residuals from the calibration data before whitening.

### 25 Apr 2026 v4

@@ 5.2 Time-Series Plot for Svalbard Temperatures in 2022, 2023, and 2024
Recall the premise that the 2022 calibration data, the 2023 training data, and the 2024 forecasting data represent realizations from the same underlying joint probability distribution; the underlying physics for corresponding months in those three years should be comparable. However, there could be aberrations perhaps caused by rapid Arctic amplification.
@@
Figure 2 shows substantial overlap across the three years. Seasonal trends are captured similarly, both in temperature values and in timing. The three time series move largely in lockstep from early to late summer, during which melting temperatures are nearly universal. There appears to be no compelling reason to reject the claim that each series consists of random variables drawn from the same joint probability distribution. In practice, the 2022 and 2024 temperatures serve as promising holdout samples for evaluating forecasts of the 2023 temperatures.
@@ 5.3 The Quantile GBM Fit of Temperature
With the quantile parameter fixed at $\tau=0.60$, the quantile gradient-boosting procedure in R ran efficiently. The shrinkage value was set to 0.0001, the interaction depth to 6, and the minimum number of observations in a terminal node to 6. These values were chosen to foster slow convergence so that the right tail of the temperature distribution would be fitted more accurately. Using 2021 data as a holdout sample  [^5] Recall that the 2021 data were used to obtain lagged values for the predictors during the earliest 2 weeks of the training data. The construction these lagged values precedes the training. to protect against overfitting indicated that approximately 27,000 iterations were appropriate for these data. [^6] With a substantially larger shrinkage value, many fewer iterations might have been adequate. However, the best shrinkage value could not be known in advance, and empirical tuning encourages “cherry picking” that can undermine later statistical inference (Kuchibhotla, Kolassa and Kuffner, 2022).
@@
Figure 3 plots the observed 2 p.m. temperatures against their fitted values. The relationship is roughly linear and positive; observed and fitted values tend to increase together. Clustering around the smoothed fitted line is somewhat tighter when the fitted values exceed the melting point, with residuals rarely larger than about $\pm 2^{\circ}\mathrm{C}$. This results in part from the larger cost for underfitting imposed by the target quantile $\tau=0.60$. There is greater data sparsity on either side of the temperature range between approximately $5^{\circ}\mathrm{C}$ and $0^{\circ}\mathrm{C}$. In short, the fit quality produced by quantile gradient boosting is encouraging. [^7] Koenker and Machado (1999) discuss measures of fit for conditional-quantile models derived from the quantile loss function. However, apparently no analogue of the multiple correlation coefficient exists for quantile regression. A suitable measure of fit is introduced later, when conformal prediction regions are presented.
@@
Overall, the fitted values track the observed temperatures rather closely and capture temperatures above $0^{\circ}\mathrm{C}$ especially well. The unusually warm days clustering in mid-July—around $10^{\circ}\mathrm{C}$—are of particular concern to climate scientists (Semenov, 2021), and they too are fitted accurately, albeit with slight underestimation. Two weeks in advance, the fitted algorithm anticipates melting temperatures effectively, including those mid-July extremes. Higher quantiles such as $Q(0.90)$ would rely on far sparser data and risk instability, making $\tau=0.60$ a reasonable balance. As with Figure 3, however, true forecasting remains to be performed.
@@ 5.5 Functional Forms
The quantile gradient-boosting algorithm learns associations between each predictor and the response, including the shapes of those relationships. The partial-dependence plots in Figure 6 show the relationships between the day counter on the left and the 14-day lagged temperature on the right with the 2 p.m. temperature response, each computed with the distributions of all other predictors unchanged. These are the two variables that dominate the fit. The black dots represent the partial-dependence values, and the solid blue lines show a loess smooth as a visual aid.
Figure 6: Partial-dependence plots for the day counter and the 14-day lagged 2 p.m. temperature. The black dots are the partial-dependence values, and the solid blue line is a loess smooth serving as a visual aid. Other predictors distributions are unchanged.
Consistent with Figure 5, strong relationships are evident. A look again at Figure 4 provides useful context here. The counter-based fitted values range from approximately $3^{\circ}\mathrm{C}$ to $3^{\circ}\mathrm{C}$, and the lagged-temperature fitted values from about $2.5^{\circ}\mathrm{C}$ to $5^{\circ}\mathrm{C}$. The first plot is roughly symmetric and concave, peaking during the summer months, consistent with the longitudinal pattern shown in Figure 4. The second plot is S-shaped, with its steepest slope beginning immediately after the melting point at $0^{\circ}\mathrm{C}$ and plateauing near $2.5^{\circ}\mathrm{C}$ at the left and $5^{\circ}\mathrm{C}$ at the right. The lagged observed temperatures have their strongest association with fitted temperatures just above the melting point. The practical significance of this threshold will be discussed in greater depth below. [^9] The short horizontal strings of dots visible in the plot represent rounding artifacts; the values differ slightly. The remaining four predictors also exhibit nonlinear relationships with the temperature response, but because of their comparatively small contributions to the fit, their partial-dependence plots are omitted in the interest of space.
@@ 6 Forecasts and Estimated Uncertainty
Addressing uncertainty in a principled way has been anticipated in the research design and algorithmic procedures employed. [^10] The material in this section is likely to be unfamiliar to many readers because it differs from traditional efforts to compute forecasting uncertainty. Its recent popularity results from far fewer assumptions that are difficult to test. Somewhat more didactic material is provided. The 2024 data serve as a pristine holdout sample of test data that can represent unlabeled cases used in genuine forecasting. [^11] The 2024 data have labels that are useful for didactic purposes and for evaluation of forecasting accuracy. But those labels have no role in the forecasts themselves and would not be known in an operational setting. Forecasts for the 2024 data are produced using the previously fitted quantile gradient-boosting algorithm and its predict() function in R. From the perspective of the trained algorithm, the 2024 lagged predictors constitute new, unlabeled data.
Obtaining the adaptive conformal prediction regions is more involved. All conformal prediction regions begin with a pre-specified coverage probability. By convention, this probability is denoted by $(1-\alpha)$, where $\alpha$ lies between 0 and 0.50. In practice, coverage probabilities range from slightly above 0.50 to nearly 1.0, with values between 0.75 and 0.95 most common. The value of $\alpha$ is determined by subjective considerations shaped by the data, the application and the certainty that subsequent decision makers will likely require. For the conformal approach used here with the Svalbard data, the coverage is $1-\alpha=0.80$. Other coverage values could be used with reasoning to follow unaffected.
@@
Exchangeability is required for valid prediction regions (Vovk, Gammerman and Shafer, 2005; Vovk et al., 2017; Angelopoulos, Barber and Bates, 2024). One key implication of exchangeability is that it does not matter for the theoretical guarantees whether a test observation is realized after all of the calibration observations are realized. Angelopoulos and Bates (2023) provide an accessible introduction to conformal prediction.
@@ 6.1 Uncertainty Results
The forecast data from 2024 provide an empirical illustration of how well the forecasting procedures perform. These data have not been used in any way to this point, and as such, share important properties with real forecasting settings that might arise in practice. However, there is additional information, ordinarily unavailable, that can be used to better document forecasting skill: the actual future temperatures being forecast and a full year of observations not just one day at a time.
@@
Figure 7: For the 2024 holdout forecast data, the mean forecasted temperature in a bin is on the horizontal axis. The percentage of cases in each of 20 equal-width bins having a true future temperature exceeding $0^{\circ}\mathrm{C}$ is on the vertical axis. Filled circles correspond to bins with mean forecasted temperature above $0^{\circ}\mathrm{C}$. Open circles correspond to those below. The solid blue line is a loess smoother.
Figure 7 provides a visual summary of these ideas using the 2024 forecast data. The horizontal axis shows the mean of the daily forecasted temperatures for each of 20 equal-width bins. Because the future temperatures are known for each day in 2024 cases, the vertical axis shows the percentage of forecasts within each bin whose true future temperature exceeds $0^{\circ}\mathrm{C}$. (The 14-day lag of each predictor ensures that these are future outcomes.) The loess overlay is S-shaped, with the greatest rate of change between about $2^{\circ}\mathrm{C}$ and $2^{\circ}\mathrm{C}$.
@@
The loess curve in Figure 7 shows a gradual transition around $0^{\circ}\mathrm{C}$ rather than an abrupt step function. The physics of melting helps explain why. Melting is a complicated process (Noto and Ulloa, 2026) and can vary for different kinds of ice (Wadhams, 2000). For example, when sea ice forms, its hexagonal lattice cannot readily incorporate salt ions. The excluded salt forms brine pockets and channels among the ice grains, leaving the ice as nearly fresh water while the brine becomes increasingly salty. As freezing continues, the concentrated brine’s freezing point can fall well below $0^{\circ}\mathrm{C}$ because the salt ions attract water molecules the lattice needs; it is more difficult for the water molecules to self-organize into the lattice structure. Unlike fresh water ice, which melts at a single fixed temperature, sea ice melts over a range of temperatures, because the brine pockets within it are already liquid at temperatures below $0^{\circ}\mathrm{C}$. As the temperature rises, progressively more of the remaining ice melts, and the brine channels widen and connect. Because a significant fraction of sea ice consists of these pre-existing liquid brine inclusions, it requires less energy to melt than an equivalent mass of fresh water ice — the brine pockets need no latent heat to liquefy, only the nearly pure ice fraction does. By the same logic, the energy released when salt water freezes is less than that released when fresh water freezes. This interpretation aligns with Figure 7.
@@ 7 Discussion
The analysis of 2 p.m. temperatures in Svalbard, Norway, combines three interlocking statistical traditions: gradient boosting with a quantile as the estimation target, an AR(1) model to construct exchangeable nonconformal scores, and adaptive conformal prediction regions computed with quantile random forests, a method robust to sparse data." To the best of the author’s knowledge, these procedures applied to three temporally separated samples has not previously appeared in the literature.
A two-week forecasting lead time before the onset of widespread surface melting is a narrow window, but it would give Arctic community administrators critical time to prepare for infrastructure resilience, public safety, and logistical challenges (Streletskiy, Shiklomanov and Christiansen, 2019; U.S. Arctic Research Commission, 2003). Local governments could limit heavy-vehicle use on thaw-sensitive roads, runways, and around pipelines; pre-position maintenance materials; and inspect culverts, bridges, and drainage channels before damage occurs (Arctic Research Consortium of the United States (ARCUS), 2021). Water managers could adjust reservoir levels or activate temporary treatment measures to mitigate siltation and contamination from meltwater inflows.
In communities dependent on ice roads, snowmobile trails, or frozen river crossings, early warnings would permit orderly resupply and fuel delivery before surface travel becomes unsafe (Chen et al., 2025). Coastal and hillside settlements could prepare for increased risks of shoreline erosion or permafrost-related landslides (The Climate Institute and The Firelight Group, 2022). Health authorities might use the lead time to issue advisories on water quality or vector-borne disease risk as standing water accumulates. [^15] Because of Arctic warming, even Iceland is now reported to host mosquitoes (Straker, 2025).
More broadly, forecasts on a two-week timescale could foster coordination among local administrators, regional governments, and research stations, improving situational awareness and resource allocation across sectors. Even modest improvements in lead time may yield substantial adaptive benefits in Arctic regions where melt conditions evolve rapidly and logistical flexibility is limited.
@@ 8 Conclusions
Projections of future global warming have long been central to IPCC assessments. The analyses presented here contribute to the complementary literature on smaller-scale local forecasting. Temperature forecasts with a two-week lead time show promise, especially when focused on Arctic thresholds associated with widespread melting. Forecasting uncertainty is explicitly addressed with adaptive conformal prediction regions, and results are presented in a manner intended to be accessible and useful to stakeholders. With thoughtfully applied algorithms and accessible data, informative short-term forecasts can be generated on an ordinary desktop computer with an internet connection.
Appendix: Statistical Narrative
@@
These linked procedures combine quantile gradient boosting, AR(1) whitening, and quantile-random-forest calibration to yield adaptive conformal prediction regions. The lower and upper bounds use the ranks of the nonconformal scores.
@@
The supervised learning algorithm is first fitted on the training data $\{(X_{t},Y_{t})\}_{t=1}^{T_{\mathrm{train}}}$. Predictions are then generated for the separate calibration data $\{X_{t}\}_{t=1}^{T}$, and their AR(1) innovations form the nonconformal scores $\{S_{t}\}_{t=1}^{T}$.
If $\{S_{t}\}_{t=1}^{T+1}$ are (approximately) exchangeable then, conditional on the fitted algorithm, the calibration scores and the new score $S_{T+1}$ are also exchangeable. Consequently,
@@
Exchangeability means that the joint distribution of the nonconformal scores is invariant to permutations of their indices. Although in practice the scores are realized in temporal order, their joint law would be unchanged had they been realized in any other order. That a new unlabeled case is realized at time T+1T{+}1 is immaterial. For the analyses of the Svalbard data, exchangeability depends on the whitening of the calibration data residuals. After whitening, the resulting innovations satisfy the exchangeability condition sufficiently well to justify the coverage guarantee above.